# Quantifying and Mitigating Privacy Risks for Tabular Generative Models

### v1

@@ 4. Empirical Analysis
In this section, we put our risk-utility quantification framework to the test on publicly available datasets that have been extensively employed in tabular data analysis and synthesis. Our code is available in the following anonymous repository: https://anonymous.4open.science/r/DP-TLDM-317C.
4.1. Datasets
@@
We employ four datasets, two small (up to 20000 samples) and two larger. Since small datasets usually make models prone to overfitting, by comparing these datasets, we can understand how dataset size and overfitting affect the quality and privacy of synthetic data. Some characteristics of the datasets are listed in Table 1.
@@
The Loan dataset (Bank, 2017) contains demographic information on 5000 customers. It holds 14 features divided into 4 different measurement categories, including binary, interval, ordinal, and nominal features. The Housing dataset (Torgo, 1990) relates to houses in a given California district and provides summary statistics based on the 1990 Census data. It comprises 20,640 instances with 1 categorical and 9 numerical features and a total of 207 missing values. The Adult dataset (Becker and Kohavi, 1996) contains information on individuals’ annual incomes and related variables. It consists of 48842 instances with 14 mixed datatype features in total, and a total of 6465 missing values. The Cardiovascular Heart Disease dataset (Dempsy, 2021) contains detailed information on the risk factors for cardiovascular disease, including 70000 instances with 13 mixed-type columns.
@@
No. Rows No. Columns Missing Values
Loan 5000 14 0
Housing 20640 10 0.1%
Adult 48842 14 0.94%
Cardio 70000 13 0
Table 1. Characteristics of four tabular datasets used in our study.
4.2. Privacy-utility Trade-off
Table 2 presents detailed results quantifying both utility and risk aspects of synthetic data for all four datasets employing the five generative models described in Section 3.1. We present the three utility metrics discussed before (i.e., resemblance, discriminability, and utility), where a higher score indicates better performance, as well as the privacy risk for the four attacks considered (Singling out, Linkability, AIA, MIA), where a lower risk indicates better performance. Due to space reasons, we keep the detailed statistics of the five MIA attacks in the appendix.
@@ 4.2. Privacy-utility Trade-off
Table 2. Quantification of Risk-Utility for Five Generative Models Across Various Datasets. Here, ”Resem.” stands for Resemblance, ”Distrim.” refers to Discriminability, ”S-out” denotes singling out attacks, and ”Link” represents linkability attacks.
@@
Regarding MIA strategies, notable effectiveness is achieved by the NaiveGroundhog (NG), HistGroundhog (HG), and Closest Distance-Hamming (CD-H) strategies, which are able to reach success rates of 60% or higher in some cases. These results are detailed in Table 4 in Appendix B.1. Remarkably, HistGroundhog consistently outperforms other MIA strategies when applied to the TabDDPM synthesizer. In contrast, the NaiveGroundhog and Closest Distance-Hamming strategies demonstrate better efficacy when employed on other synthesizers.
@@ 5. DP-TLDM
Figure 2. The latent tabular diffusion model. Given the original tabular data with both continuous and categorical features, the autoencoder first encodes both features into a cohesive latent space, with the protection of Differential Privacy (DP). The Latent Diffusion Model then executes a Gaussian diffusion process within the latent space.
The essential background of diffusion models for tabular data is introduced in Section 5.1. Following this, in Section 5.2, we delineate the motivation behind the development of our latent tabular diffusion model, as well as details regarding the two components. Furthermore, the inclusion of DP-enhanced training and Differential Privacy measures are presented in Sections 5.4 and 5.3, respectively.
5.1. Diffusion Primer
@@
In the realm of modeling tabular data, the inherent heterogeneity among features necessitates tailored approaches for accurate modeling. TabDDPM (Kotelnikov et al., 2023) addresses this challenge by adopting different methods for noising and denoising continuous and categorical features, as shown in Figure 3. TabDDPM employs Gaussian diffusion following (Ho et al., 2020), where the forward process gradually adds Gaussian noise to the input data, which eventually ends at $p_{con}(x_{T})=\mathcal{N}(x_{T};\mathbf{0},\mathbf{I})$. Conversely, in the reverse process, a neural network is trained to predict the added noise, thereby facilitating the denoising of the data.
Figure 3. The architecture of TabDDPM where different methods are adopted for continuous and discrete features separately. Continuous features undergo normalization and are handled by the Gaussian diffusion process, whereas categorical features are one-hot encoded and diffused using the Multinomial diffusion process.
Meanwhile, categorical features are handled using Multinomial diffusion, as proposed by (Hoogeboom et al., 2021), with each categorical feature having a distinct Multinomial diffusion process. For a categorical feature with $K$ classes, during the forward process, uniform noise over the $K$ classes is applied to corrupt the one-hot encoded categorical feature, ending at the categorical distribution $p_{dis}(x_{T})=\mathcal{C}(x_{T};1/K)$. Subsequently, the reverse process leverages a neural network to predict the probability vector to recover the noised data.
@@
5.2. Tabular Latent Diffusion Model (TLDM)
@@
To address the mentioned limitations, we propose the latent tabular diffusion model following (Rombach et al., 2022). In our latent tabular diffusion model, both the continuous and categorical features are transferred to a unified continuous latent space by training an autoencoder. Subsequently, a unified diffusion model is deployed to noise and denoise the continuous latent features. The decoder component of the autoencoder is then employed to convert the denoised latent representation back to the original features.
@@ 5.2.1. Autoencoder
To train the autoencoder, we follow the common setting in Variational AutoEncoders (Kingma and Welling, 2013), minimizing as loss function the negative Evidence Lower-Bound (ELBO), defined as
@@
5.2.2. Latent Diffusion Model
@@
(2) $q(z^{t}|z^{0})=\mathcal{N}(z^{t};\sqrt{\bar{\alpha}^{t}}z^{0},(1-\bar{\alpha}^{t})\mathbf{I}),$
@@
To learn the reverse denoising transitions, we adopt the methodology proposed in (Ho et al., 2020). The crux of this approach involves estimating the added noise. Thus, the training objective of the diffusion component is formulated as minimizing the loss
@@
5.3. Differential Privacy Framework
To introduce a privacy protection in the latent space we employ the $f$-DP framework (Dong et al., 2021) since it is capable to provide better bounds on the privacy leakage under composition, which is key in the training of neural models, which is done iteratively by means of stochastic gradient descent. These better bounds result in a more faithful privacy-utility tradeoff analysis.
@@ 5.3. Differential Privacy Framework
In our paper, we adopt the $f$-DP framework to elevate privacy protection. This approach offers a clearer and more intuitive privacy explanation, encapsulating all necessary details to derive established DP metrics. Moreover, $f$-DP achieves a tighter privacy bound than traditional $(\varepsilon,\delta)$-DP, allowing for a more precise privacy evaluation (Dong et al., 2021; Wang et al., 2024).
Differential Privacy (DP), as introduced by Dwork et al. (Dwork et al., 2006), is a foundational framework for preserving the privacy of individuals’ data within datasets. It quantifies the impact of an individual’s data on the output of a randomized algorithm, ensuring minimal influence and thus protecting privacy.
@@
Transitioning from traditional $(\varepsilon,\delta)$-DP analysis, the $f$-DP framework, proposed in (Dong et al., 2021), offers a refined perspective that relies on framing the adversary’s challenge as a hypothesis testing problem. This framework introduces a trade-off function $f$ that represents the trade-off between false negatives (FN) and false positives (FP) in distinguishing between datasets $D$ and $D^{\prime}$.
@@
Thanks to its functional definition, the $f$-DP framework is able to provide much tighter composition bounds than other existing definitions of DP. In fact, $f$-DP encompasses $(\epsilon,\delta)$-DP as a special case, wherein a mechanism is $(\epsilon,\delta)$-DP if and only if it conforms to $f_{\epsilon,\delta}$-DP, with $f_{\epsilon,\delta}(\alpha)=\max\{0,1-\delta-e^{\epsilon}\alpha,(1-\delta-\alpha)e^{-\epsilon}\}$.
@@
The Differentially Private Stochastic Gradient Descent algorithm (DP-SGD) (Abadi et al., 2016) was designed for the differentially private training of neural networks. It achieves differential privacy by individually clipping (IC) the gradient of each individual sample within each mini-batch and adding Gaussian noise $\mathcal{N}(0,(C\sigma)^{2}\textbf{I})$ as
(3) $\tilde{g}_{r}\leftarrow\frac{1}{|B|}({\textstyle\sum_{i\in B}[g_{r}(x_{i})]_{C}+\mathcal{N}(0,(C\sigma)^{2}\textbf{I})}).$
@@
Let $N$ be the dataset size, $b=\mathbb{E}[|B|]$ the sample (mini batch) size, and $\sigma$ the standard deviation of the Gaussian noise used in DP-SGD 555 We consider probabilistic sampling as in the Opacus library (Yousefpour et al., 2021) and use noise parameter $C\cdot\sigma$ and normalize with $C$ rather than $2C$. as in (3). Thus, $N/b$ equals the number of rounds in a single epoch and letting $E$ denote the total number of epochs, the total number of rounds is $R=(N/b)\cdot E$.
Then DP-SGD is $C_{b/N}(G_{\sigma^{-1}})^{\otimes R}$-DP where $C_{b/N}$ is an operator representing the effect of subsampling in DP-SGD, $G_{\sigma^{-1}}$ is a Gaussian trade-off function characterizing the differential privacy (called Gaussian DP) due to adding Gaussian noise in DP-SGD, and the operator $\otimes T$ describes composition (of privacy leakage) over $R$ rounds.
Following the asymptotic analysis in (Dong et al., 2021), DP-SGD converges to a $\mu$-Gaussian DP defined as
@@
The ideal trade-off function is defined as $f(\alpha)=1-\alpha$, representing random guessing by the adversary; hence, it implies no privacy leakage. Since optimal trade-off functions are symmetric around the diagonal, the separation between $1-\alpha$ and $G_{\mu}(\alpha)$ can be measured as the Euclidean distance between the point $(\frac{1}{2},\frac{1}{2})$ on the curve $1-\alpha$ and the point $(a,a)$ on the curve $G_{\mu}(\alpha)$, i.e., where $G_{\mu}(a)=a$. Here, $G_{\mu}(a)=\Phi(\Phi^{-1}(1-a)-\mu)$, with $\Phi(\cdot)$ being the cumulative distribution function of the standard normal distribution. Thus the separation is denoted as
(4) $sep=\sqrt{2}\left|a-\frac{1}{2}\right|,\quad\text{s.t. }G_{\mu}(a)=a.$
@@
Figure 4. The separation between the ideal curve and the trade-off function.
@@
5.4. Two-stage DP-SGD Training
Instead of using the traditional individual clipping as in (3), as pointed out in (Nguyen et al., 2023), a better way is to utilize batch clipping (BC) for DP training, i.e.,
(5) $\tilde{g}_{r}\leftarrow{\textstyle\left[\frac{1}{|B|}\sum_{i\in B}g_{r}(x_{i})\right]_{C}+\mathcal{N}(0,(C\sigma)^{2}\textbf{I})},$
@@ 5.4. Two-stage DP-SGD Training
Second, batch clipping enables training Batch Normalization Layers in neural networks with a robust DP guarantee. As highlighted in (Nguyen et al., 2023), current implementations for IC that use batch normalization on the extensive training dataset lead to correlations among the updates across training rounds. Since these correlations are not considered, IC does not yield a solid DP guarantee from a theoretical perspective. However, batch normalization with BC over corresponding mini-batches can provide a more solid DP argument within the flimit-$f-$DP framework.
@@
1:Input: Tabular data $X=\{x_{1},\ldots,x_{N}\}$ with $N$ samples, training epochs $E_{1}$ and $E_{2}$, batch sizes $B_{1}$ and $B_{2}$, noise scale $\sigma$, gradient norm bound $C$, and total timestep $T$.
2:Output: Trained encoder $\mathcal{E}$, decoder $\mathcal{D}$, and diffusion noise estimation network $\epsilon_{\theta}$.
3:Initialize encoder $\mathcal{E}$, decoder $\mathcal{D}$, and diffusion noise estimation network $\epsilon_{\theta}$
4:for $e_{1}=1$ to $E_{1}$ do $\triangleright$ Train the Autoencoder component
@@
6:         Compute and average batch gradients:
7:         $\bar{g}_{r_{1_{\mathcal{E}}}}\leftarrow\frac{1}{B_{1}}\nabla_{\mathcal{E}}\mathcal{L}_{AE}(\mathcal{E},X_{r_{1}})$
8:         $\bar{g}_{r_{1_{\mathcal{D}}}}\leftarrow\frac{1}{B_{1}}\nabla_{\mathcal{D}}\mathcal{L}_{AE}(\mathcal{D},X_{r_{1}})$
9:         Clip batch gradients:
10:         $\left[\bar{g}_{r_{1_{\mathcal{E}}}}\right]_{C}\leftarrow\bar{g}_{r_{1_{\mathcal{E}}}}/\max(1,\|\bar{g}_{r_{1_{\mathcal{E}}}}\|_{2}/C)$
11:         $\left[\bar{g}_{r_{1_{\mathcal{D}}}}\right]_{C}\leftarrow\bar{g}_{r_{1_{\mathcal{D}}}}/\max(1,\|\bar{g}_{r_{1_{\mathcal{D}}}}\|_{2}/C)$
12:         Add noise:
13:         $\tilde{g}_{r_{1_{\mathcal{E}}}}\leftarrow\left[\bar{g}_{r_{1_{\mathcal{E}}}}\right]_{C}+\mathcal{N}(0,(C\sigma)^{2}\mathbf{I})$
14:         $\tilde{g}_{r_{1_{\mathcal{D}}}}\leftarrow\left[\bar{g}_{r_{1_{\mathcal{D}}}}\right]_{C}+\mathcal{N}(0,(C\sigma)^{2}\mathbf{I})$
15:         Gradient descent on $\mathcal{E}$ and $\mathcal{D}$ with $\tilde{g}_{r_{1_{\mathcal{E}}}}$ and $\tilde{g}_{r_{1_{\mathcal{D}}}}$
@@
18:
19:$Z^{0}=\mathcal{E}(X)$triangleright Generate the latent space from the DP-trained Autoencoder
20:for $e_{2}=1$ to $E_{2}$ do $\triangleright$ Train the Latent Diffusion component
21:     for $r_{2}=1$ to $\lceil N/B_{2}\rceil$ do
22:         $Z_{r_{2}}^{0}\sim q(Z^{0})$
23:         Sample time step and true noise:
24:         $t\sim\text{Uniform}(\{1,\ldots,T\})$, $\epsilon\sim\mathcal{N}(0,\mathbf{I})$
25:         Compute gradient: $g_{\epsilon_{\theta}}\leftarrow\nabla_{\epsilon_{\theta}}\mathcal{L}_{DF}(\epsilon,\epsilon_{\theta},t,Z_{r_{2}}^{0})$
26:         Gradient descent on $\epsilon_{\theta}$ with $g_{\epsilon_{\theta}}$
27:     end for
28:end for
6. Evaluation Analysis
In this section, we evaluate the proposed DP-TLDM on the aforementioned four datasets, employing the same quality and privacy risk metrics used in Section 4. We aim to answer if DP-TLDM can take advantage of the privacy protection from the DP mechanism without degrading the synthetic data quality. We specifically compare DP-TLDM against two other baselines, DP-CTGAN and DP-TabDDPM, by applying the DP-SGD training algorithm on CTGAN and TabDDPM, which represent the state-of-the-art GAN and diffusion-based generative models, respectively.

### v2

@@ 3 Empirical Analysis
In this section, we put our risk-utility quantification framework described in Appendix B.3 to the test on publicly available datasets that have been extensively employed in tabular data analysis and synthesis.
3.1 Datasets
@@
We employ four datasets, two small (up to 20000 samples) and two larger. Since small datasets usually make models prone to overfitting, by comparing these datasets, we can understand how dataset size and overfitting affect the quality and privacy of synthetic data. Some characteristics of the datasets are listed in Table 4 in the appendix.
@@
The Loan dataset [3] contains demographic information on 5000 customers. It holds 14 features divided into 4 different measurement categories, including binary, interval, ordinal, and nominal features. The Housing dataset [43] relates to houses in a given California district and provides summary statistics based on the 1990 Census data. It comprises 20,640 instances with 1 categorical and 9 numerical features and a total of 207 missing values. The Adult dataset [4] contains information on individuals’ annual incomes and related variables. It consists of 48842 instances with 14 mixed datatype features in total, and a total of 6465 missing values. The Cardiovascular Heart Disease dataset [11] contains detailed information on the risk factors for cardiovascular disease, including 70000 instances with 13 mixed-type columns.
@@
3.2 Privacy-utility Trade-off
Table 3 presents detailed results quantifying both utility and risk aspects of synthetic data for all four datasets employing the five generative models described in Appendix B.1. We present the three utility metrics discussed before (i.e., resemblance, discriminability, and utility), where a higher score indicates better performance, as well as the privacy risk for the four attacks considered (Singling out, Linkability, AIA, MIA), where a lower risk indicates better performance. Due to space reasons, we keep the detailed statistics of the five MIA attacks in the appendix.
@@ 3.2 Privacy-utility Trade-off
Table 1: Quantification of Risk-Utility for Five Generative Models Across Various Datasets. Here, “Resem.” stands for Resemblance,“Distrim.” refers to Discriminability, “S-out” denotes singling out attacks, and ”Link” represents linkability attacks.
@@
Regarding MIA strategies, notable effectiveness is achieved by the NaiveGroundhog (NG), HistGroundhog (HG), and Closest Distance-Hamming (CD-H) strategies, which are able to reach success rates of 60% or higher in some cases. These results are detailed in Table 5 in Appendix D.1. Remarkably, HistGroundhog consistently outperforms other MIA strategies when applied to the TabDDPM synthesizer. In contrast, the NaiveGroundhog and Closest Distance-Hamming strategies demonstrate better efficacy when employed on other synthesizers.
@@ 4 DP-TLDM
Figure 1: The latent tabular diffusion model. Given the original tabular data with both continuous and categorical features, the autoencoder first encodes both features into a cohesive latent space, with the protection of Differential Privacy (DP). The Latent Diffusion Model then executes a Gaussian diffusion process within the latent space.
The essential background of diffusion models for tabular data is introduced in Section 4.1. Following this, in Section 4.2, we delineate the motivation behind the development of our latent tabular diffusion model, as well as details regarding the two components. Furthermore, the inclusion of DP-enhanced training and Differential Privacy measures are presented in Sections 4.4 and 4.3, respectively.
4.1 Diffusion Primer
@@
In the realm of modeling tabular data, the inherent heterogeneity among features necessitates tailored approaches for accurate modeling. TabDDPM [30] addresses this challenge by adopting different methods for noising and denoising continuous and categorical features, as shown in Figure 2. TabDDPM employs Gaussian diffusion following [22], where the forward process gradually adds Gaussian noise to the input data, which eventually ends at $p_{con}(x_{T})=\mathcal{N}(x_{T};\mathbf{0},\mathbf{I})$. Conversely, in the reverse process, a neural network is trained to predict the added noise, thereby facilitating the denoising of the data.
Figure 2: The architecture of TabDDPM where different methods are adopted for continuous and discrete features separately. Continuous features are handled by the Gaussian diffusion process, whereas categorical features are one-hot encoded and diffused using the Multinomial diffusion process.
Meanwhile, categorical features are handled using Multinomial diffusion, as proposed by [23], with each categorical feature having a distinct Multinomial diffusion process. For a categorical feature with $K$ classes, during the forward process, uniform noise over the $K$ classes is applied to corrupt the one-hot encoded categorical feature, ending at the categorical distribution $p_{dis}(x_{T})=\mathcal{C}(x_{T};1/K)$. Subsequently, the reverse process leverages a neural network to predict the probability vector to recover the noised data.
@@
4.2 Tabular Latent Diffusion Model (TLDM)
@@
To address the mentioned limitations, we propose the latent tabular diffusion model following [37]. In our latent tabular diffusion model, both the continuous and categorical features are transferred to a unified continuous latent space by training an autoencoder. Subsequently, a unified diffusion model is deployed to noise and denoise the continuous latent features. The decoder component of the autoencoder is then employed to convert the denoised latent representation back to the original features.
@@ 4.2.1 Autoencoder
To train the autoencoder, we follow the common setting in Variational AutoEncoders [29], minimizing as loss function the negative Evidence Lower-Bound (ELBO), defined as
@@
4.2.2 Latent Diffusion Model
@@
$q(z^{t}|z^{0})=\mathcal{N}(z^{t};\sqrt{\bar{\alpha}^{t}}z^{0},(1-\bar{\alpha}^{t})\mathbf{I}),$ (1)
@@
To learn the reverse denoising transitions, we adopt the methodology proposed in [22]. The crux of this approach involves estimating the added noise. Thus, the training objective of the diffusion component is formulated as minimizing the loss
@@
4.3 Differential Privacy Framework
To introduce a privacy protection in the latent space we employ the $f$-DP framework [13] since it is capable to provide better bounds on the privacy leakage under composition, which is key in the training of neural models, which is done iteratively by means of stochastic gradient descent. These better bounds result in a more faithful privacy-utility tradeoff analysis.
@@ 4.3 Differential Privacy Framework
In our paper, we adopt the $f$-DP framework to elevate privacy protection. This approach offers a clearer and more intuitive privacy explanation, encapsulating all necessary details to derive established DP metrics. Moreover, $f$-DP achieves a tighter privacy bound than traditional $(\varepsilon,\delta)$-DP, allowing for a more precise privacy evaluation [13, 45].
Differential Privacy (DP), as introduced by Dwork et al. [15], is a foundational framework for preserving the privacy of individuals’ data within datasets. It quantifies the impact of an individual’s data on the output of a randomized algorithm, ensuring minimal influence and thus protecting privacy.
@@
Transitioning from traditional $(\varepsilon,\delta)$-DP analysis, the $f$-DP framework, proposed in [13], offers a refined perspective that relies on framing the adversary’s challenge as a hypothesis testing problem. This framework introduces a trade-off function $f$ that represents the trade-off between false negatives (FN) and false positives (FP) in distinguishing between datasets $D$ and $D^{\prime}$.
@@
Thanks to its functional definition, the $f$-DP framework can provide much tighter composition bounds than other existing definitions of DP. $f$-DP encompasses $(\epsilon,\delta)$-DP as a special case, wherein a mechanism is $(\epsilon,\delta)$-DP if and only if it conforms to $f_{\epsilon,\delta}$-DP, with $f_{\epsilon,\delta}(\alpha)=\max\{0,1-\delta-e^{\epsilon}\alpha,(1-\delta-\alpha)e^{-\epsilon}\}$.
@@
The Differentially Private Stochastic Gradient Descent algorithm (DP-SGD) [1] was designed for the differentially private training of neural networks. It achieves differential privacy by individually clipping (IC) the gradient of each individual sample within each mini-batch and adding Gaussian noise $\mathcal{N}(0,(C\sigma)^{2}\textbf{I})$ as to the gradient
$\tilde{g}_{r}\leftarrow\frac{1}{|B|}({\textstyle\sum_{i\in B}[g_{r}(x_{i})]_{C}+\mathcal{N}(0,(C\sigma)^{2}\textbf{I})}).$ (2)
@@
Let $N$ be the dataset size, $b=\mathbb{E}[|B|]$ the sample (mini batch) size, and $\sigma$ the standard deviation of the Gaussian noise used in DP-SGD 111 We consider probabilistic sampling as in the Opacus library [50] and use noise parameter $C\cdot\sigma$ and normalize with $C$ rather than $2C$. as in (2). Thus, $N/b$ equals the number of rounds in a single epoch and letting $E$ denote the total number of epochs, the total number of rounds is $R=(N/b)\cdot E$.
Then DP-SGD is $C_{b/N}(G_{\sigma^{-1}})^{\otimes R}$-DP where $C_{b/N}$ is an operator representing the effect of subsampling in DP-SGD, $G_{\sigma^{-1}}$ is a Gaussian trade-off function characterizing the differential privacy (called Gaussian DP) due to adding Gaussian noise in DP-SGD, and the operator $\otimes R$ describes composition (of privacy leakage) over $R$ rounds.
Following the asymptotic analysis in [13], DP-SGD converges to a $\mu$-Gaussian DP defined as
@@
The ideal trade-off function is defined as $f(\alpha)=1-\alpha$, representing random guessing by the adversary; hence, it implies no privacy leakage. Since optimal trade-off functions are symmetric around the diagonal, separation between $1-\alpha$ and $G_{\mu}(\alpha)$ can be measured as Euclidean distance between the point $(\frac{1}{2},\frac{1}{2})$ on the curve $1-\alpha$ and the point $(a,a)$ on the curve $G_{\mu}(\alpha)$, i.e., where $G_{\mu}(a)=a$. Here, $G_{\mu}(a)=\Phi(\Phi^{-1}(1-a)-\mu)$, with $\Phi(\cdot)$ being the cumulative distribution function of the standard normal distribution. Thus the separation is denoted as
$sep=\sqrt{2}\left|a-\frac{1}{2}\right|,\quad\text{s.t. }G_{\mu}(a)=a.$ (3)
@@
Figure 3: The separation between the ideal curve and the trade-off function.
@@
4.4 Two-stage DP-SGD Training
Instead of using the traditional individual clipping (IC) as in (2), as pointed out in [35], a better way is to utilize batch clipping (BC) for DP training, i.e.,
$\tilde{g}_{r}\leftarrow{\textstyle\left[\frac{1}{|B|}\sum_{i\in B}g_{r}(x_{i})\right]_{C}+\mathcal{N}(0,(C\sigma)^{2}\textbf{I})},$ (4)
@@ 4.4 Two-stage DP-SGD Training
1:Input: Tabular data $X=\{x_{1},\ldots,x_{N}\}$, epochs $E_{1}$, $E_{2}$, batch sizes $B_{1}$, $B_{2}$, noise scale $\sigma$, norm bound $C$, timestep $T$
2:Output: Encoder $\mathcal{E}$, decoder $\mathcal{D}$, noise network $\epsilon_{\theta}$
3:Initialize $\mathcal{E}$, $\mathcal{D}$, $\epsilon_{\theta}$
4:for $e_{1}=1$ to $E_{1}$ do
@@
6:         Compute $\bar{g}_{r_{1_{\mathcal{E}}}}\leftarrow\frac{1}{B_{1}}\nabla_{\mathcal{E}}\mathcal{L}_{AE}(\mathcal{E},X_{r_{1}})$, $\bar{g}_{r_{1_{\mathcal{D}}}}\leftarrow\frac{1}{B_{1}}\nabla_{\mathcal{D}}\mathcal{L}_{AE}(\mathcal{D},X_{r_{1}})$
7:         Clip and add noise for $\mathcal{E}$: $\tilde{g}_{r_{1_{\mathcal{E}}}}\leftarrow\frac{\bar{g}_{r_{1_{\mathcal{E}}}}}{\max(1,\|\bar{g}_{r_{1_{\mathcal{E}}}}\|_{2}/C)}+\mathcal{N}(0,(C\sigma)^{2}\mathbf{I})$
8:         Clip and add noise for $\mathcal{D}$: $\tilde{g}_{r_{1_{\mathcal{D}}}}\leftarrow\frac{\bar{g}_{r_{1_{\mathcal{D}}}}}{\max(1,\|\bar{g}_{r_{1_{\mathcal{D}}}}\|_{2}/C)}+\mathcal{N}(0,(C\sigma)^{2}\mathbf{I})$
9:     end for
10:end for
11:$Z^{0}=\mathcal{E}(X)$
12:for $e_{2}=1$ to $E_{2}$ do
13:     for $r_{2}=1$ to $\lceil N/B_{2}\rceil$ do
14:         Sample $Z_{r_{2}}^{0}\sim q(Z^{0})$, $t\sim\text{Uniform}(\{1,\ldots,T\})$, $\epsilon\sim\mathcal{N}(0,\mathbf{I})$
15:         Compute $g_{\epsilon_{\theta}}\leftarrow\nabla_{\epsilon_{\theta}}\mathcal{L}_{DF}(\epsilon,\epsilon_{\theta},t,Z_{r_{2}}^{0})$ and update $\epsilon_{\theta}$
@@
Second, batch clipping enables training Batch Normalization Layers in neural networks with a robust DP guarantee. As highlighted in [35], current implementations for IC that use batch normalization on the extensive training dataset lead to correlations among the updates across training rounds. Since these correlations are not considered, IC does not yield a solid DP guarantee from a theoretical perspective. However, batch normalization with BC over corresponding mini-batches can provide a solid DP argument within the flimit-$f-$DP framework.
5 Performance Evaluation
In this section, we evaluate the proposed DP-TLDM on the aforementioned four datasets, employing the same quality and privacy risk metrics used in Section 0.C. We aim to answer if DP-TLDM can take advantage of the privacy protection from the DP mechanism without degrading the synthetic data quality. We specifically compare DP-TLDM against two other baselines, DP-CTGAN and DP-TabDDPM, by applying the DP-SGD training algorithm on CTGAN and TabDDPM, which represent the state-of-the-art GAN and diffusion-based generative models, respectively.